# Benign overfitting in leaky ReLU networks with moderate input dimension

### v1

@@ 4 Approximate margin maximization and generalization: insight from linear models
In this section we outline proofs for the analogues of Theorems 3.2 and 3.3 in the context of linear models. The arguments are thematically similar and clearer to present. We provide complete proofs of benign and non-benign overfitting for linear models in Appendix C.
@@
Let ${\bm{N}}\in\mathbb{R}^{n\times d}$ denote a random matrix whose rows are drawn mutually i.i.d. from $\mathcal{N}(\bm{0}_{d},\tfrac{1}{d}({\bm{I}}_{d}-{\bm{v}}{\bm{v}}^{T}))$. If $d=\Omega\left(n+\log\frac{1}{\delta}\right)$, then with probability at least $1-\delta$,
$c\leq\sigma_{n}({\bm{N}})\leq\sigma_{1}({\bm{N}})\leq C.$
@@ 1. Generalization bounds based on the SNR:
A little analysis yields that $X\sim\mathcal{N}\left(\sqrt{\gamma},\tfrac{1-\gamma}{d}\|{\bm{z}}\|^{2}\right)$ and the desired upper bound therefore follows from Hoeffding’s inequality,
@@
$\mathbb{P}(X\leq 0)\geq\frac{1}{2}-\sqrt{\frac{d\gamma}{2\pi(1-\gamma)}}\frac{a_{v}}{\|{\bm{z}}\|}.$
@@ 2. Upper bound the norm of the max-margin classifier:
In order to use the approximate max-margin property we require an upper bound on $\|{\bm{w}}^{*}\|$. As by definition $\|{\bm{w}}^{*}\|\leq\|\tilde{{\bm{w}}}\|$, it suffices to construct a vector $\tilde{{\bm{w}}}$ which interpolates the data and has small norm. Using that the noise matrix of the data is well-conditioned with high probability, we achieve this by strategically constructing the signal and noise components of $\tilde{{\bm{w}}}$. This yields the bound
@@
where the arguments of the max function originate from a small and large $\gamma$ regime respectively.
@@ 3. Lower bound the SNR using the approximate margin maximization property:
To prove non-benign overfitting, we compute an upper bound for the ratio $a_{v}/\|{\bm{z}}\|$ rather than a lower bound. Since the model perfectly fits the training data with margin one,
@@
for all $i\in[n]$. If $\gamma$ is small and $|\sqrt{\gamma}a_{v}|$ is large, then $\|{\bm{w}}\|$ will be large, contradicting the approximate margin maximization. Hence the above inequality implies that $\sqrt{1-\gamma}\hat{y}_{i}\langle{\bm{n}}_{i},{\bm{z}}\rangle$ is large for all $i\in[n]$. Since the noise is well-conditioned, this can only happen when $\|{\bm{z}}\|$ is large. This gives us a lower bound on $\|{\bm{z}}\|$. As before, we can also upper bound $a_{v}$ by $\|{\bm{w}}\|$, giving us an upper bound on the SNR $a_{v}/\|{\bm{z}}\|$. By the second generalization bound in step 1, the classifier generalizes poorly and exhibits non-benign overfitting.
@@ 4.1 From linear models to leaky ReLU networks
where $a_{j}\in\mathbb{R}$ and ${\bm{z}}_{j}\perp{\bm{v}}$. In the linear case ±$\pm a_{v}$ can be interpreted as the activation of the linear classifier on ±$\pm{\bm{v}}$ respectively: in terms of magnitude the signal activation is the same in either case and thus we measure the alignment of the linear model with the signal using $|a_{v}|$. For leaky ReLU networks we define their activation on ±$\pm{\bm{v}}$ respectively as $A_{1}=f({\bm{W}},{\bm{v}})$ and $A_{-1}=f({\bm{W}},-{\bm{v}})$, and then define the alignment of the network as $A_{\min}=\min\{A_{1},A_{-1}\}$. Considering the alignment of the network with the noise, then if ${\bm{Z}}\in\mathbb{R}^{2m\times d}$ denotes a matrix whose $j$-th row is ${\bm{z}}_{j}$, then we measure the alignment of the network using $\|{\bm{Z}}\|_{F}$. As a result, analogous to $a_{v}/\|{\bm{z}}\|$, the key ratio from a generalization perspective in the context of a leaky ReLU network is $A_{\min}/\|{\bm{Z}}\|_{F}$. The proof Theorems 3.2 and 3.3 then follow the same outline as Steps 1-3 above but with additional non-trivial technicalities.
@@ 5 Conclusion
In this work we have proven conditions under which leaky ReLU networks trained on binary classification tasks exhibit benign and non-benign overfitting. We have substantially relaxed the necessary assumptions on the input data compared with prior work; instead of requiring nearly orthogonal data with $d=\Omega(n^{2}\log n)$ or higher, we only need $d=\Omega(n)$. We achieve this by using the distribution of singular values of the noise rather than specific correlations between noise vectors. Our emphasis was on networks trained by gradient descent with the hinge loss, but we establish a new framework that is general enough to accommodate any algorithm which is approximately margin maximizing. A natural question for future work is for which other loss functions and algorithms the approximate margin maximization condition will hold, and whether the model will exhibit benign overfitting in these regimes. Another interesting question is on the role of linear separability and whether there is an analogue of margin maximization for non-linearly separable data which leads to benign overfitting.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1928930 and by the Alfred P. Sloan Foundation under grant G-2021-16778, while the authors EG and DN were in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Fall 2023 semester. EG and DN were also partially supported by NSF DMS 2011140. EG was also supported by a NSF Graduate Research Fellowship under grant DGE 2034835. GM and KK were partly supported by NSF CAREER DMS 2145630 and DFG SPP 2298 Theoretical Foundations of Deep Learning grant 464109215. GM was also partly supported by NSF grant CCF 2212520, ERC Starting Grant 757983 (DLT), and BMBF in DAAD project 57616814 (SECAI).
@@ Appendix A Preliminaries on random vectors
Let ${\bm{n}}\sim\mathcal{N}(\bm{0}_{d},d^{-1}({\bm{I}}_{d}-{\bm{v}}{\bm{v}}^{T}))$ and suppose ${\bm{z}}\in\operatorname{span}(\{{\bm{v}}\})^{\perp}$. There exists a $C>0$ such that with probability at least $1-\delta$
@@ Appendix B Upper bounding the norm of the max-margin classifier of the data
which establishes property 4. ∎
@@ C.1 Sufficient conditions for benign and harmful overfitting
$\mathbb{P}(y\langle{\bm{w}},{\bm{x}}\rangle\leq 0)\geq\frac{1}{2}-\sqrt{\frac{d\gamma}{2\pi(1-\gamma)}}\frac{a_{v}}{\|{\bm{z}}\|}.$
@@
$\geq\frac{1}{2}-\frac{1}{\sqrt{2\pi}}\left(\sqrt{\frac{\gamma d}{1-\gamma}}\frac{a_{v}}{\|{\bm{z}}\|}\right).\qed$
@@
In the context of the data model described in Section 2, let ${\bm{w}}^{\ast}$ be a max-margin linear classifier of the training data. Let $\mathcal{A}:\mathbb{R}^{n\times d}\times\{\pm 1\}^{n}\rightarrow\mathbb{R}^{d}$ be a learning algorithm which is approximately margin-maximizing, Definition 2.3. For $\delta,\epsilon\in(0,1]$, let $d=\Omega\left(|\mathcal{A}|^{2}n\log\frac{1}{\epsilon}+\log\frac{1}{\delta}\right)$. Then with probability at least $1-\delta$ over the randomness of the training data $({\bm{X}},\hat{{\bm{y}}})$, the following hold.
@@
    If $\gamma=\Omega(|\mathcal{A}|^{2}n^{-1})$ and $k=O(|\mathcal{A}|^{-2}n)$, then the generalization error of $\mathcal{A}({\bm{X}},\hat{{\bm{y}}})$ is upper bounded by $\epsilon$.
@@
    If $\gamma=O(|\mathcal{A}|^{-2}\epsilon^{2}d^{-1})$, then the generalization error of $\mathcal{A}({\bm{X}},\hat{{\bm{y}}})$ is lower bounded by $\frac{1}{2}-\epsilon$.
@@
First we establish (A). For algorithms learning linear models then by definition $|\mathcal{A}\|\geq 1$ as $\|{\bm{w}}\|\leq\|{\bm{w}}^{*}\|$. Therefore the given condition on $d$ above ensures $d=\Omega\left(n+\log\frac{1}{\delta}\right)$, and therefore Lemmas 4.1 and B.3 show that with probability at least $1-\frac{\delta}{2}$, $\|{\bm{N}}\|^{2},\|{\bm{N}}_{\mathcal{G}}\|^{2}\leq C$ and $\|{\bm{w}}^{*}\|\leq C\sqrt{\frac{1}{\gamma}+\frac{k}{1-\gamma}}$. Here ${\bm{N}}_{\mathcal{G}}$ denotes the matrix formed by taking only the rows of ${\bm{N}}$ which satisfy $\beta=1$. We denote this event $\omega$ and condition on it in all that follows for the proof of (A). As $\mathcal{A}$ is approximately max margin then given equation 8 we have for all $i\in\mathcal{G}$
@@
which further implies $n\leq\frac{C|\mathcal{A}|^{2}}{\gamma}+C|\mathcal{A}|^{2}k$ for some other constant $C$. For this inequality to hold, either $\frac{C|\mathcal{A}|^{2}}{\gamma}\geq\frac{n}{2}$ or $C|\mathcal{A}|^{2}k\geq\frac{n}{2}$. With $k\leq\frac{n}{4C|\mathcal{A}|^{2}}$ and $\gamma\geq\frac{4C|\mathcal{A}|^{2}}{n}$, neither of these can be true and therefore we conclude that $\sqrt{\gamma}a_{v}\geq\frac{1}{2}$. Then
@@
$\geq\frac{C}{|\mathcal{A}|^{2}\gamma}\frac{1}{\frac{1}{\gamma}+\frac{n}{1-\gamma}}$
where the positive constant $C$ may vary between inequalities. Letting $({\bm{x}},y)$ denote a test point pair, then by Lemma C.1 with $\gamma\geq n^{-1}$ and $d\geq 2C^{-1}|\mathcal{A}|^{2}n\log(1/\epsilon)$ it follows that
@@
$\leq\exp\left(-\frac{Cd}{|\mathcal{A}|^{2}\left(\frac{1-\gamma}{\gamma}+n\right)}\right)$
$\leq\exp\left(-\frac{Cd}{|\mathcal{A}|^{2}\left(\frac{1}{\gamma}+n\right)}\right)$
$\leq\exp\left(-\frac{Cd}{2|\mathcal{A}|^{2}n}\right)$
$\leq\epsilon.$
Hence the generalization error is at most $\epsilon$ when $\omega$ occurs, which happens with probability at least $1-\delta$. This establishes (A).
We now turn our attention to (B). As justified in the proof of part (A) we can set $d=\Omega\left(n+\log\frac{1}{\delta}\right)$. Therefore from Lemmas 4.1 and B.4 with probability at least $1-\delta$ it holds that $\|{\bm{N}}_{\mathcal{G}}\|^{2}\leq C$ and $\|{\bm{w}}^{*}\|\leq\frac{C\sqrt{n}}{\sqrt{1-\gamma}}$. We denote this event $\omega^{\prime}$ and condition on it in all that follows for the proof of (B). In particular,
@@
where the constant $C>0$ may vary between inequalities. Letting $({\bm{x}},y)$ denote a test point pair, by Lemma C.1 with $\gamma\leq\frac{\epsilon^{2}}{Cd|\mathcal{A}|^{2}}$ it follows that
@@
$\geq\frac{1}{2}-\sqrt{Cd\gamma|\mathcal{A}|^{2}}$
$\geq\frac{1}{2}-\epsilon.$
Hence the generalization error is at least $\frac{1}{2}-\epsilon$ when $\omega^{\prime}$ occurs, which happens with probability at least $1-\delta$. This establishes (B). ∎

### v2

@@ 4 Approximate margin maximization and generalization: insight from linear models
In this section we outline proofs for the analogues of Theorems 3.2 and 3.4 in the context of linear models. The arguments are thematically similar and clearer to present. We provide complete proofs of benign and non-benign overfitting for linear models in Appendix C.
@@
Let ${\bm{N}}\in\mathbb{R}^{n\times d}$ denote a random matrix whose rows are drawn mutually i.i.d. from $\mathcal{N}(\bm{0}_{d},\tfrac{1}{d}({\bm{I}}_{d}-{\bm{v}}{\bm{v}}^{T}))$. If $d=\Omega\left(n+\log\frac{1}{\delta}\right)$, then there exists constants $C_{1}$ and $C_{2}$ such that, with probability at least $1-\delta$,
$C_{1}\leq\sigma_{n}({\bm{N}})\leq\sigma_{1}({\bm{N}})\leq C_{2}.$
@@ 1. Generalization bounds based on the SNR:
As the noise is normally distributed, $X\sim\mathcal{N}\left(\sqrt{\gamma},\tfrac{1-\gamma}{d}\|{\bm{z}}\|^{2}\right)$ and the desired upper bound therefore follows from Hoeffding’s inequality,
@@
$\mathbb{P}(y\langle{\bm{w}},{\bm{x}}\rangle\leq 0)\geq\max\left\{\frac{1}{2}-\sqrt{\frac{d\gamma}{2\pi(1-\gamma)}}\frac{a_{v}}{\|{\bm{z}}\|},\frac{1}{4}\exp\left(-\frac{6d}{\pi}\frac{\gamma}{1-\gamma}\frac{a_{v}^{2}}{\|{\bm{z}}\|^{2}}\right)\right\}.$
@@ 2. Upper bound the norm of the max-margin classifier:
In order to use the approximate max-margin property we require an upper bound on $\|{\bm{w}}^{*}\|$. As by definition $\|{\bm{w}}^{*}\|\leq\|\tilde{{\bm{w}}}\|$, it suffices to construct a vector $\tilde{{\bm{w}}}$ that interpolates the data and has small norm. Using that the noise matrix of the data is well-conditioned with high probability, we achieve this by strategically constructing the signal and noise components of $\tilde{{\bm{w}}}$. This yields the bound
@@
where the arguments of the min function originate from a small and large $\gamma$ regime respectively.
@@ 3. Lower bound the SNR using the approximate margin maximization property:
4. Upper bound the SNR using the zero loss condition:
For the generalization lower bound, we compute an upper bound for the ratio $a_{v}/\|{\bm{z}}\|$ rather than a lower bound. Since the model perfectly fits the training data with margin one,
@@ 4. Upper bound the SNR using the zero loss condition:
for all $i\in[n]$. The above inequality implies that $\sqrt{1-\gamma}\hat{y}_{i}\langle{\bm{n}}_{i},{\bm{z}}\rangle$ is at least $\sqrt{\gamma}a_{v}$ for all corrupt points. Since the noise is well-conditioned, this gives a lower bound on $\|{\bm{z}}\|$ in terms of $a_{v}$ and hence an upper bound on the SNR $a_{v}/\|{\bm{z}}\|$. By the second generalization lower bound in step 1, the generalization error is bounded below at a similar exponential rate to the upper bound.
5. Upper bound the SNR using the zero loss condition and maximum margin property:
To prove non-benign overfitting, we again compute an upper bound for the ratio $a_{v}/\|{\bm{z}}\|$. We return to the zero loss condition:
for all $i\in[n]$. If $\gamma$ is small and $|\sqrt{\gamma}a_{v}|$ is large, then $\|{\bm{w}}\|$ will be large, contradicting the approximate margin maximization. Hence the above inequality implies that $\sqrt{1-\gamma}\hat{y}_{i}\langle{\bm{n}}_{i},{\bm{z}}\rangle$ is large for all $i\in[n]$. Since the noise is well-conditioned, this can only happen when $\|{\bm{z}}\|$ is large. This gives us a lower bound on $\|{\bm{z}}\|$. As before, we can also upper bound $a_{v}$ by $\|{\bm{w}}\|$, giving us an upper bound on the SNR $a_{v}/\|{\bm{z}}\|$. By the first generalization lower bound in step 1, the classifier generalizes poorly and exhibits non-benign overfitting.
@@ 4.1 From linear models to leaky ReLU networks
where $a_{j}\in\mathbb{R}$ and ${\bm{z}}_{j}\perp{\bm{v}}$. In the linear case ±$\pm a_{v}$ can be interpreted as the activation of the linear classifier on ±$\pm{\bm{v}}$ respectively: in terms of magnitude the signal activation is the same in either case and thus we measure the alignment of the linear model with the signal using $|a_{v}|$. For leaky ReLU networks we define their activation on ±$\pm{\bm{v}}$ respectively as $A_{1}=f({\bm{W}},{\bm{v}})$ and $A_{-1}=-f({\bm{W}},-{\bm{v}})$,
and then define the alignment of the network as $A_{\min}=\min\{A_{1},A_{-1}\}$. Considering the alignment of the network with the noise, then if ${\bm{Z}}\in\mathbb{R}^{2m\times d}$ denotes a matrix whose $j$-th row is ${\bm{z}}_{j}$, then we measure the alignment of the network using $\|{\bm{Z}}\|_{F}$. As a result, analogous to $a_{v}/\|{\bm{z}}\|$, the key ratio from a generalization perspective in the context of a leaky ReLU network is $A_{\min}/\|{\bm{Z}}\|_{F}$. The proof Theorems 3.2, 3.3, and 3.4 then follow the same outline as Steps 1-3 above but with additional non-trivial technicalities.
@@ 5 Conclusion
In this work we have proven conditions under which leaky ReLU networks trained on binary classification tasks exhibit benign and non-benign overfitting. We have substantially relaxed the necessary assumptions on the input data compared with prior work; instead of requiring nearly orthogonal data with $d=\Omega(n^{2}\log n)$ or higher, we only need $d=\Omega(n)$. We achieve this by using the distribution of singular values of the noise rather than specific correlations between noise vectors. Our emphasis was on networks trained by gradient descent with the hinge loss, but we establish a new framework that is general enough to accommodate any algorithm that is approximately margin maximizing.
There are a few limitations of our results which would be natural questions to address in future work. While we improve upon existing results in our dependence on the input dimension of the data, we still require that the training dataset is linearly separable. This leaves open the question of whether an overparameterized network will perfectly fit the training data and generalize well for lower dimensional data, or satisfy a similar margin maximization condition. We also focus mainly on two-layer networks with fixed outer layer weights trained with the hinge loss. It would be interesting to investigate whether analogous results hold for deeper architectures or different loss functions and data models.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. DMS-1928930 and by the Alfred P. Sloan Foundation under grant G-2021-16778, while the authors EG and DN were in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Fall 2023 semester.
EG and DN were also partially supported by NSF DMS 2011140. EG was also supported by a NSF Graduate Research Fellowship under grant DGE 2034835.
GM and KK were partly supported by NSF CAREER DMS 2145630 and DFG SPP 2298 Theoretical Foundations of Deep Learning grant 464109215. GM was also partly supported by NSF grant CCF 2212520, ERC Starting Grant 757983 (DLT), and BMBF in DAAD project 57616814 (SECAI).
@@ Appendix A Preliminaries on random vectors
Let ${\bm{n}}\sim\mathcal{N}(\bm{0}_{d},d^{-1}({\bm{I}}_{d}-{\bm{v}}{\bm{v}}^{T}))$ and suppose ${\bm{z}}\in\operatorname{span}(\{{\bm{v}}\})^{\perp}$.
There exists a $C>0$ such that with probability at least $1-\delta$
@@ Appendix B Upper bounding the norm of the max-margin classifier of the data
which establishes property 4.
@@ C.1 Sufficient conditions for benign and harmful overfitting
$\mathbb{P}(y\langle{\bm{w}},{\bm{x}}\rangle\leq 0)\geq\max\left\{\frac{1}{2}-\sqrt{\frac{d\gamma}{2\pi(1-\gamma)}}\frac{a_{v}}{\|{\bm{z}}\|},\frac{1}{4}\exp\left(-\frac{6d}{\pi}\frac{\gamma}{1-\gamma}\frac{a_{v}^{2}}{\|{\bm{z}}\|^{2}}\right)\right\}.$
@@
$\geq\frac{1}{2}-\frac{1}{\sqrt{2\pi}}\left(\sqrt{\frac{\gamma d}{1-\gamma}}\frac{a_{v}}{\|{\bm{z}}\|}\right).$
Another bound can be obtained using the following inequality (polya1949remarks, (1.5)):
[…]
We proceed
[…]
$\geq\frac{1}{2}-\frac{1}{2}\sqrt{1-\exp\left(-\frac{2\gamma da_{v}^{2}}{\pi(1-\gamma)\|{\bm{z}}\|^{2}}\right)}$
$\geq\frac{1}{2}-\frac{1}{2}\left(1-\frac{1}{2}\exp\left(-\frac{2\gamma da_{v}^{2}}{\pi(1-\gamma)\|{\bm{z}}\|^{2}}\right)\right)$
$\geq\frac{1}{4}\exp\left(-\frac{2\gamma da_{v}^{2}}{\pi(1-\gamma)\|{\bm{z}}\|^{2}}\right).\qed$
@@
In the context of the data model described in Section 2, let ${\bm{w}}^{\ast}$ be a max-margin linear classifier of the training data. Let $\mathcal{A}:\mathbb{R}^{n\times d}\times\{\pm 1\}^{n}\rightarrow\mathbb{R}^{d}$ be a learning algorithm which is approximately margin-maximizing, Definition 2.3. For $\delta\in(0,1]$, let $d=\Omega\left(n+\log\frac{1}{\delta}\right)$. Then with probability at least $1-\delta$ over the randomness of the training data $({\bm{X}},\hat{{\bm{y}}})$, the following hold with $\epsilon$ denoting the generalization error of $\mathcal{A}({\bm{X}},\hat{{\bm{y}}})$.
@@
    If $\gamma=\Omega(|\mathcal{A}|^{2}n^{-1})$ and $k=O(|\mathcal{A}|^{-2}n)$, then $\exp\left(-C_{1}dk^{-1}\right)\leq\epsilon\leq\exp\left(-C_{2}dk^{-1}|\mathcal{A}|^{-2}\right)$ for fixed positive constants $C_{1}$ and $C_{2}$.
@@
    If $\gamma=O(|\mathcal{A}|^{-2}d^{-1})$, then $\epsilon\geq\frac{1}{2}-\sqrt{Cd\gamma|\mathcal{A}|^{2}}$ for a fixed positive constant $C$.
@@
First we establish (A). As $d=\Omega\left(n+\log\frac{1}{\delta}\right)$, Lemmas 4.1 and B.3 show that with probability at least $1-\frac{\delta}{2}$, $\|{\bm{N}}\|^{2},\|{\bm{N}}_{\mathcal{G}}\|^{2},\|{\bm{N}}_{\mathcal{B}}\|\leq C$ and $\|{\bm{w}}^{*}\|\leq C\sqrt{\frac{1}{\gamma}+\frac{k}{1-\gamma}}$. Here ${\bm{N}}_{\mathcal{G}}$ and ${\bm{N}}_{\mathcal{B}}$ denote the matrices formed by taking only the rows of ${\bm{N}}$ which satisfy $\beta=1$ and $\beta=-1$, respectively. We denote this event $\omega$ and condition on it in all that follows for the proof of (A).
As $\mathcal{A}$ is approximately max margin then given equation 8 we have for all $i\in\mathcal{G}$
@@
which further implies $n\leq\frac{C|\mathcal{A}|^{2}}{\gamma}+C|\mathcal{A}|^{2}k$ for some other constant $C$. For this inequality to hold, either $\frac{C|\mathcal{A}|^{2}}{\gamma}\geq\frac{n}{2}$ or $C|\mathcal{A}|^{2}k\geq\frac{n}{2}$. With $k\leq\frac{n}{4C|\mathcal{A}|^{2}}$ and $\gamma\geq\frac{1}{k}$, neither of these can be true and therefore we conclude that $\sqrt{\gamma}a_{v}\geq\frac{1}{2}$. Then
@@
$\geq\frac{C}{2|\mathcal{A}|^{2}\gamma}\frac{1-\gamma}{k}$
for a positive constant $C$. Letting $({\bm{x}},y)$ denote a test point pair, then by Lemma C.1 it follows that for a different constant $C$
@@
$\leq\exp\left(-\frac{Cd}{|\mathcal{A}|^{2}k}\right).$
Hence the generalization error is at most $\epsilon$ when $\omega$ occurs, which happens with probability at least $1-\delta$. This establishes the upper bound of (A).
For the lower bound of (A), since ${\bm{w}}$ is a linear classifier,
$\langle{\bm{n}}_{i},{\bm{z}}\rangle\geq\frac{1}{\sqrt{1-\gamma}}(1+a_{v}\sqrt{\gamma})\geq a_{v}\sqrt{\frac{\gamma}{1-\gamma}}$
for all $i\in\mathcal{B}$, from which we conclude $|\langle{\bm{n}}_{i},{\bm{z}}\rangle|\geq a_{v}\sqrt{\gamma}$. This implies
$\|{\bm{N}}_{\mathcal{B}}{\bm{z}}\|\geq a_{V}\sqrt{\frac{k\gamma}{1-\gamma}}$
Along with $\|{\bm{N}}_{\mathcal{B}}{\bm{z}}\|\leq\|{\bm{N}}_{\mathcal{B}}\|\|{\bm{z}}\|\leq C\|{\bm{z}}\|$ we conclude
$\|{\bm{z}}\|\geq\frac{a_{v}}{C}\sqrt{\frac{k\gamma}{1-\gamma}}.$
With this bound we then bound
$\frac{a_{v}}{\|{\bm{z}}\|}$ $\leq C\cdot\sqrt{\frac{1-\gamma}{k\gamma}}.$
By Lemma C.1 we then can bound
$\mathbb{P}(y\langle{\bm{w}},{\bm{x}}\rangle\leq 0\;|\;\omega)$ $\geq\frac{1}{4}\exp\left(-\frac{6d}{\pi}\frac{\gamma}{1-\gamma}\frac{a_{v}^{2}}{\|{\bm{z}}\|^{2}}\right)$
$\geq\exp\left(-\frac{Cd}{k}\right)$
for a new constant $C$, provided $a_{v}$ is positive. In the last line we can bound $\frac{d}{k}$ below as $d\geq n$ and $k=O(n)$. If $a_{v}$ is negative then the generalization error is at least $\frac{1}{2}$, which is still bounded below by $\exp(-Cd/k)$.
We now turn our attention to (B). As $d=\Omega(n+\log\tfrac{1}{\delta})$, from Lemmas 4.1 and B.4 with probability at least $1-\delta$ it holds that $\|{\bm{N}}_{\mathcal{G}}\|^{2}\leq C$ and $\|{\bm{w}}^{*}\|\leq\frac{C\sqrt{n}}{\sqrt{1-\gamma}}$. We denote this event $\omega^{\prime}$ and condition on it in all that follows for the proof of (B). In particular,
@@
where the constant $C>0$ may vary between inequalities. Letting $({\bm{x}},y)$ denote a test point pair, by Lemma C.1 it follows that
@@
$\geq\frac{1}{2}-\sqrt{Cd\gamma|\mathcal{A}|^{2}}.$
Hence the generalization error is at least $\frac{1}{2}-\sqrt{Cd\gamma|\mathcal{A}|^{2}}$ when $\omega^{\prime}$ occurs, which happens with probability at least $1-\delta$. This establishes (B). ∎